Hello,

I need some help solving this problem:

Al's flower garden measures 25 feet by 35 feet. He has enough crushed stone to cover 325 square feet. If he uses the stone to build a path around his garden, what is the maximum possible width of the path.

Thank you

Answer by scott8148(6628)

let x=(width of path)

area of garden with path is (25+2x)(35+2x) ... (25)(35)+120x+4x^2
area of garden is (25)(35)
the difference in the areas is the area of the path ... 325

4x^2+120x=325 ... 4x^2+120x-325=0 ... (4x+130)(x-2.5)=0